OpenStudy (anonymous):

Solve each system using substitution. 2=2y-x 23=5y-4x

OpenStudy (mathstudent55):

Solve the first equation for x: 2 = 2y - x Add x to both sides: x + 2 = 2y Subtract 2 from both sides: x = 2y - 2 Now substitute what x equals in second equation. 23 = 5y - 4x 23 = 5y - 4(2y - 2) Simplify the right side: 23 = 5y - 8y + 8 23 = -3y + 8 Subtract 8 from both sides: 15 = -3y Divide both sides by -3: -5 = y Now that you have y = -5, substitute this value of y in first equation: 2 = 2y - x 2 = 2(-5) - x 2 = -10 - x 12 = -x x = -12 Answer: x = -12, y = -5

OpenStudy (anonymous):

ty :D

OpenStudy (mathstudent55):

wlcm